if the area of a rectangle is 52 and perimeter is 34 and length=width*2+5 what is length and width
Answer by Alan3354(69443)

if the area of a rectangle is 52 and perimeter is 34 and length=width*2+5 what is length and width
------------
Area = L*W = 52
Perimeter = 2L + 2W = 34 --> L + W = 17
-----------
Try pairs of factors of 52 (there aren't many).
Find a pair with a sum of 17.